# Realization of the Tellegen Effect in Resonant Optical Metasurfaces

### v1

@@ Abstract
The nonreciprocal magnetoelectric effect in isotropic Tellegen materials enables exotic phenomena such as axion-modified electrodynamics [1, 2, 3, 4, 5, 6, 7, 8] and fosters the development of magnet-free nonreciprocal media [9, 10, 11]. Although predicted over 75 years ago [12] and observed in only a handful of natural materials with very low strength [13], the strong optical Tellegen effect has remained experimentally elusive. Here, we report the first experimental demonstration of a resonant optical isotropic Tellegen effect in a metasurface, showcasing a response that is 100 times greater than that of any known natural material. This optical metasurface, consisting of randomly distributed cobalt-silicon nanoscatterers with strong shape anisotropy, utilizes spontaneous magnetization to achieve a robust Tellegen effect without the need for an external magnetic field. In addition to the Tellegen response, the metasurface exhibits both gyroelectric and gyromagnetic effects [14], contributing to nonreciprocal cross-polarized light reflection. We introduce a technique to independently extract the amplitudes of these three effects using conventional magneto-optical single-side-illumination measurements. The observation of the resonant Tellegen effects in the optical frequency range may lead to the experimental observation of axionic electrodynamics and compact bias-free nonreciprocal optical devices.
@@
The isotropic Tellegen effect is expected to have a significant impact on both fundamental and applied physics [21]. It should be clearly distinguished from its anisotropic counterpart, the magnetochiral effect [22, 23], as it necessitates a diagonal magnetoelectric tensor with a non-zero trace. Seminal works have demonstrated that the electrodynamic equations governing isotropic Tellegen materials take the same form as those describing a hypothetical axion medium within the framework of quantum field theory [1, 2]. The axion is an elementary particle theorized to address two major issues in modern physics: the strong charge conjugation and parity (CP) problem [24] and the existence of dark matter [25]. Tellegen-type nonreciprocal materials have attracted considerable interest as potential enablers of exotic concepts of axion electrodynamics in practical materials, such as dyon quasiparticles, anyon statistics, the Witten effect, and axionic polaritons [2, 26, 27, 28]. They may also play a crucial role in the search for axions [6, 7, 5]. Moreover, Tellegen materials promise a range of exotic phenomena, especially in the optical regime, including nonreciprocal light reflection [29, 10] and transmission [9], photonic topological edge states [30], PT-symmetric magnetoelectric energy density [31], and spin-dependent thermal radiation [32].
While the isotropic Tellegen effect could be measured at microwave frequencies for certain multiferroic materials [21, 33, 34], it already becomes two orders of magnitude smaller than the refractive index in the terahertz range [35] and five orders of magnitude smaller in the optical regime [36, 37]. This drastic reduction at optical frequencies is attributed to the fact that both cyclotron and Larmor frequencies of electrons responsible for breaking the Lorentz reciprocity typically lie in the gigahertz range. Thus, artificial Tellegen composites designs have been confined to the microwave frequencies: from individual constituents (meta-atoms) [38, 39, 40] to the two-dimensional arrays (metasurfaces) [41, 42, 11]. The quest for optical Tellegen effect has inspired three recent conceptual proposals for Tellegen metamaterials. The first two approaches, based on rapid temporal modulation of material properties [43, 44] and multilayer antiferromagnetic configuration [8, 45], are not achievable with the current manufacturing technology. The third theoretical proposal based on metasurfaces with spontaneous magnetization in magnetic two-component nanoparticles has not been experimentally validated either [10].
@@
The isotropic Tellegen effect is primarily an interface phenomenon manifesting itself as a nonreciprocal cross-polarized light reflection at a vacuum-matter interface. For spatially uniform slabs, the effect remains undetectable in light transmission [29]. Unlike magneto-optical effects of both gyroelectric and gyromagnetic types, Tellegen media exhibit a direction-independent Kerr effect due to isotropy and do not exhibit the Faraday effect [10]. This distinction makes measuring the Tellegen effect more challenging, as its observable manifestation—the “isotropic” Kerr effect—depends solely on the concentration of Tellegen meta-atoms rather than their total number or the slab thickness [29, p. 261]. Therefore, the fabrication and experimental verification of high-density optical Tellegen metasurfaces, where interface phenomena are readily observable, represent the most straightforward approach to exploring the Tellegen effect.
@@
Figure 1: Optical Tellegen metasurface design. (a) Illustration of the metasurface with optimized optical Tellegen meta-atoms consisting of cobalt (top) and silicon (bottom) parts. The metasurface is illuminated from opposite directions by normally incident light (a slight oblique incidence is depicted for visual clarity). The gradient color indicates that the reflected and transmitted light contain both co- and cross-polarized components. The inset depicts a close-up image of the meta-atom with close to single-domain vertical magnetization ${\bf M\mit}_{\rm s}$ in the cobalt. It also shows the high-frequency magnetic field distribution inside the nanocone when illuminated by a standing wave with an electric field antinode positioned at the center of the meta-atom. The field distribution is calculated at a free-space wavelength of 830 nm and reveals an $x$-oriented magnetic dipole moment ${\bf m\mit}$ induced solely by the incident $x$-polarized electric field through the isotropic (diagonal) Tellegen effect. (b)–(e) Simulated complex polarizabilities of the metasurface in (a), normalized by the free-space impedance $\eta_{0}$.
@@
We select the Tellegen meta-atom as a nanoscale cone-shaped scatterer, shown in the inset of Fig. 1(a). The upper part of the nanocone is cobalt, with a height $h_{\rm Co}=119$ nm and a base diameter $D_{\rm Co}=76$ nm. The lower part is amorphous silicon, with a height $h_{\rm Si}=20$ nm and a base diameter $D_{\rm Si}=89$ nm. This geometry has an apex angle of approximately $35^{\circ}$, enabling straightforward fabrication via bottom-up hole-mask colloidal lithography (HCL). Cobalt possesses one of the largest magneto-optical parameters in the visible spectrum [48]. The optimal base diameter of the cobalt part, $D_{\rm Co}$, was determined by fabricating test metasurfaces consisting of cobalt-only nanocones to achieve a configuration closest to the ideal single-domain state (see Supplementary Section 1) that arises due to the strong shape anisotropy [49] and remains unaffected by the addition of the silicon. That is, the cobalt part can sustain strong vertical remanent magnetization close to the saturation magnetization $M_{\rm s}$ in the absence of an external magnetic field. The optimal height of the silicon part was determined by full-wave simulations to induce a magnetic-type Mie resonance [10], thereby enhancing the Tellegen response of the meta-atom (see Supplementary Section 2). The magnetic field distribution of this resonant mode inside the nanocone is depicted in the inset of Fig. 1(a) at a free-space wavelength of 830 nm. Here, the excitation consists of a standing wave with an $x$-polarized electric field antinode located at the center of the meta-atom. This field alone induces an $x$-oriented magnetic dipole moment ${\bf m\mit}$ inside the nanocone – a distinctive signature of the Tellegen effect [16].
@@
We first consider the metasurface in free space with meta-atoms in a square periodic lattice (Fig. 1(a).) We analyze the metasurface response using full-wave simulations and extract the collective polarizabilities, which relate the induced electric and magnetic dipole moments ($\mathbf{p}$ and $\mathbf{m}$) to the incident electric and magnetic fields ($\mathbf{E}$ and $\mathbf{H}$[16]:
$\begin{array}[]{c}\displaystyle{\bf p\mit}=\left[\alpha_{\rm e}\overline{\overline{I}}+\alpha_{\rm ge}\overline{\overline{J}}\right]\cdot{\bf E\mit}+\left[\alpha_{\rm te}\overline{\overline{I}}+\alpha_{\rm om}\overline{\overline{J}}\right]\cdot{\bf H\mit},\vspace{6pt}\\ \displaystyle{\bf m\mit}=\left[\alpha_{\rm m}\overline{\overline{I}}+\alpha_{\rm gm}\overline{\overline{J}}\right]\cdot{\bf H\mit}+\left[\alpha_{\rm te}\overline{\overline{I}}+\alpha_{\rm om}\overline{\overline{J}}\right]\cdot{\bf E\mit},\end{array}$ (1)
where $\overline{\overline{I}}$ represents the two-dimensional identity matrix, and $\overline{\overline{J}}$ denotes the antisymmetric (skew-symmetric) two-dimensional matrix characterized by the nonzero components $J_{yx}=-J_{xy}=1$. Light is incident on the metasurface along the $z$-axis (normal to the surface), which is also the uniaxial direction of the metasurface. The collective polarizabilities $\alpha_{\rm e}$ and $\alpha_{\rm m}$ represent the ordinary reciprocal electric and magnetic responses, respectively. The terms $\alpha_{\rm ge}$ and $\alpha_{\rm gm}$ correspond to the nonreciprocal gyroelectric and gyromagnetic responses, while $\alpha_{\rm te}$ and $\alpha_{\rm om}$ describe the nonreciprocal Tellegen and reciprocal omega responses, respectively. All the polarizabilities are defined with respect to the base of the nanocone.
@@
By evaluating the full scattering matrix of the metasurface through full-wave simulations, we compute the three nonreciprocal collective polarizabilities (see Supplementary Section 3), as shown in Figs. 1(b)–(d). From Fig. 1(e), it can be observed that all three polarizabilities exhibit similar magnitudes, necessitating illumination from both sides for their accurate extraction. At approximately 830 nm, the Tellegen response exhibits a Lorentzian-type resonance. Additionally, the Tellegen meta-atom demonstrates a resonant gyromagnetic response, whose presence at optical frequencies was only recently observed in metamaterials [46].
@@
Experimental extraction of the Tellegen effect
@@
where we introduce $C_{\rm ge}$, $C_{\rm gm}$, and $C_{\rm te}$ as the weight amplitudes of the gyroelectric, gyromagnetic, and Tellegen effects in the cross-polarized reflected light. These amplitudes depend on the free-space wavelength $\lambda$, periodicity of the metasurface $a$, as well as material properties and thicknesses of the dielectric spacer and metallic layer. Importantly, these coefficients are independent of the meta-atoms and represent only the environment surrounding the metasurface. The free-space impedance $\eta_{0}$ is included in (3) to ensure that the polarizabilities remain normalized and have consistent dimensions.
@@
Generally, in our method of Tellegen effect extraction using three metasurfaces, one selects three arbitrary spacer thicknesses $h_{1}$, $h_{2}$, $h_{3}$ to obtain a system of three linearly independent equations for $R_{\rm cr}^{(1)}$, $R_{\rm cr}^{(2)}$, and $R_{\rm cr}^{(3)}$ as in (3), which can then be solved with respect to the unknown $\alpha_{\rm ge}$, $\alpha_{\rm gm}$, and $\alpha_{\rm te}$. The complex cross-polarized reflection coefficients are determined experimentally by performing conventional magneto-optical polar Kerr measurements [50, 51, 52]. Specifically, one can write for the $(j)$-th metasurface $R_{\rm cr}^{(j)}=(\theta_{j}+i\epsilon_{j})R_{\rm co}^{(j)}$, where $\theta$ and $\epsilon$ are the Kerr rotation and ellipticity, respectively, and $R_{\rm co}$ is the co-polarized complex reflection coefficient [14]. Without loss of generality, here we choose the dielectric spacer thicknesses as $h_{1}=0$ nm, $h_{2}=60$ nm, and $h_{3}=120$ nm approximately corresponding to the maxima of the three weight amplitudes within the wavelength range of interest (around 830 nm).
To determine the Tellegen effect of the proposed optical metasurface, we follow the schematics depicted in Fig. 2. First, we fabricate three samples with aforementioned spacer thicknesses (Figs. 2(a)–(c)). Using the magneto-optical measurements, we obtain three sets of the Kerr rotation and ellipticity $\theta_{j}$, $\epsilon_{j}$. Next, we find complex-valued reflection coefficients $R_{\rm cr}^{(j)}$ and weight amlpitudes $C_{\rm ge}$, $C_{\rm gm}$, and $C_{\rm te}$. Finally, by solving the system of three linear equations, we retrieve the unknown polarizabilities $\alpha_{\rm ge}$, $\alpha_{\rm gm}$, and $\alpha_{\rm te}$ (Supplementary Section 4).
Experimental realization of the optical Tellegen metasurface
Figure 3: Fabricated Tellegen metasurface and its characterization. (a)–(b) SEM images of one of the fabricated Tellegen metasurfaces with dielectric spacer thickness of $h=60$ nm (top and tilted views). The inset in (b) presents a zoomed-in image of a single meta-atom, artificially colored to visualize the cobalt and silicon parts with the averaged geometrical parameters of the nanocone used in full-wave simulations: $h_{\rm Si}=20$ nm, $h_{\rm Co}=104$ nm, $D_{\rm Si}=90$ nm, and $D_{\rm Co}=84$ nm. The averaged half-apex angle of the nanocone is $85^{\circ}$. (c) Schematic illustration of the polar magneto-optic Kerr effect (MOKE) experimental setup. Detailed information about the setup is available in the Methods section. (d)–(f) Experimental and simulated magneto-optical Kerr spectra for the three fabricated metasurfaces under an external magnetic field fixed at $B=1.3$ T. Here, the positive-value magnetic field denotes the direction from the base of the meta-atom to its apex. The simulated spectra correspond to periodic metasurfaces with geometric parameters averaged from the fabricated samples. (g)–(i) The gyroelectric, gyromagnetic, and Tellegen polarizabilities extracted from the data in (d)–(f).

### v2

@@ Abstract
The nonreciprocal magnetoelectric effect in Tellegen materials enables exotic phenomena such as axion-modified electrodynamics and fosters the development of magnet-free nonreciprocal media. As the nonreciprocal counterpart to the well-known chiral electromagnetic response, it offers a parallel framework in which many concepts developed for chiral materials can be translated to Tellegen media, potentially unlocking new avenues for fundamental studies and applications. Although predicted over 75 years ago and observed in only a handful of natural materials with very low strength, the strong optical Tellegen effect has remained experimentally elusive. Here, we report the first experimental demonstration of a resonant optical diagonal Tellegen effect in a metasurface, showcasing a response that is 100 times greater than that of any known natural material. This optical metasurface, consisting of randomly distributed cobalt-silicon nanoscatterers with strong shape anisotropy, utilizes spontaneous magnetization to achieve a robust Tellegen effect without the need for an external magnetic field. In addition to the Tellegen response, the metasurface exhibits both gyroelectric and gyromagnetic effects, contributing to nonreciprocal cross-polarized light reflection. We introduce a technique to independently extract the amplitudes of these three effects using conventional magneto-optical single-side-illumination measurements. The observation of the resonant Tellegen effects in the optical frequency range may lead to the experimental observation of axionic electrodynamics and compact bias-free nonreciprocal optical devices.
I Introduction
@@
The diagonal Tellegen effect is expected to have a significant impact on both fundamental and applied physics [14]. It should be clearly distinguished from its off-diagonal counterpart, the magnetochiral effect [45, 59], as it necessitates a diagonal magnetoelectric tensor with a non-zero trace. Seminal works have demonstrated that the electrodynamic equations governing isotropic Tellegen materials take the same form as those describing a hypothetical axion medium within the framework of quantum field theory [53, 64]. The axion is an elementary particle theorized to address two major issues in modern physics: the strong charge conjugation and parity (CP) problem [37] and the existence of dark matter [1]. Tellegen-type nonreciprocal materials have attracted considerable interest as potential enablers of exotic concepts of axion electrodynamics in practical materials, such as dyon quasiparticles, anyon statistics, the Witten effect, and axionic polaritons [64, 26, 57, 6]. They may also play a crucial role in the search for axions [30, 34, 13]. Moreover, Tellegen materials promise a range of exotic phenomena, especially in the optical regime, including nonreciprocal light reflection [28, 47] and transmission [39], photonic topological edge states [17], PT-symmetric magnetoelectric energy density [8], and spin-dependent thermal radiation [22].
While the diagonal Tellegen effect could be measured at microwave frequencies for certain multiferroic materials [14, 12, 41], it already becomes two orders of magnitude smaller than the refractive index in the terahertz range [25] and five orders of magnitude smaller in the optical regime [46, 24]. This drastic reduction at optical frequencies is attributed to the fact that both cyclotron and Larmor frequencies of electrons responsible for breaking the Lorentz reciprocity typically lie in the gigahertz range. Thus, artificial Tellegen composites designs have been confined to the microwave frequencies: from individual constituents (meta-atoms) [20, 60, 31] to the two-dimensional arrays (metasurfaces) [21, 44, 66]. The quest for optical Tellegen effect has inspired three recent conceptual proposals for Tellegen metamaterials. The first two approaches, based on rapid temporal modulation of material properties [40, 51] and multilayer antiferromagnetic configuration [52, 48], are not achievable with the current manufacturing technology. The third theoretical proposal based on metasurfaces with spontaneous magnetization in magnetic two-component nanoparticles has not been experimentally validated either [47].
@@ II.1 Metasurface design
The isotropic Tellegen effect is manifested by a nonreciprocal cross-polarized light reflection at a vacuum-matter interface. For spatially uniform slabs, the effect remains undetectable in light transmission [28]. Unlike magneto-optical effects of both gyroelectric and gyromagnetic types, Tellegen media exhibit a direction-independent Kerr effect due to isotropy and do not exhibit the Faraday effect [47]. This distinction makes measuring the Tellegen effect more challenging, as its observable manifestation—the “isotropic” Kerr effect—depends solely on the concentration of Tellegen meta-atoms rather than their total number or the slab thickness [28, p. 261]. Therefore, the fabrication and experimental verification of high-density optical Tellegen metasurfaces, where interface phenomena are readily observable, represent the most straightforward approach to exploring the Tellegen effect.
@@
Figure 1: Optical Tellegen metasurface design. (a) Illustration of the metasurface with optimized optical Tellegen meta-atoms consisting of cobalt (top) and silicon (bottom) parts. The metasurface is illuminated from opposite directions by normally incident light (a slight oblique incidence is depicted for visual clarity). The gradient color indicates that the reflected and transmitted light contain both co- and cross-polarized components. The inset depicts a close-up image of the meta-atom with close to single-domain vertical magnetization ${\bf M}_{\rm s}$ in the cobalt. It also shows the high-frequency magnetic field distribution inside the nanocone when illuminated by a standing wave with an electric field antinode positioned at the center of the meta-atom. The field distribution is calculated at a free-space wavelength of 830 nm and reveals an $x$-oriented magnetic dipole moment $\bf m$} induced solely by the incident $x$-polarized electric field through the isotropic (diagonal) Tellegen effect. (b)–(e) Simulated complex polarizabilities of the metasurface in (a), normalized by the free-space impedance $\eta_{0}$.
@@
We select the Tellegen meta-atom as a nanoscale cone-shaped scatterer, shown in the inset of Fig. 1(a). The upper part of the nanocone is cobalt, with a height $h_{\rm Co}=119$ nm and a base diameter $D_{\rm Co}=76$ nm. The lower part is amorphous silicon, with a height $h_{\rm Si}=20$ nm and a base diameter $D_{\rm Si}=89$ nm. This geometry has an apex angle of approximately $35^{\circ}$, enabling straightforward fabrication via bottom-up hole-mask colloidal lithography (HCL). Cobalt possesses one of the largest magneto-optical parameters in the visible spectrum [62]. The optimal base diameter of the cobalt part, $D_{\rm Co}$, was determined by fabricating test metasurfaces consisting of cobalt-only nanocones to achieve a configuration closest to the ideal single-domain state (see Supplementary Section 1) that arises due to the strong shape anisotropy [11] and remains unaffected by the addition of the silicon. That is, the cobalt part can sustain strong vertical remanent magnetization close to the saturation magnetization $M_{\rm s}$ in the absence of an external magnetic field. The optimal height of the silicon part was determined by full-wave simulations to induce a magnetic-type Mie resonance [47], thereby enhancing the Tellegen response of the meta-atom (see Supplementary Section 2). The magnetic field distribution of this resonant mode inside the nanocone is depicted in the inset of Fig. 1(a) at a free-space wavelength of 830 nm. Here, the excitation consists of a standing wave with an $x$-polarized electric field antinode located at the center of the meta-atom. This field alone induces an $x$-oriented magnetic dipole moment $\mathbf{m}$ inside the nanocone. This behavior is a direct manifestation of the Tellegen effect [4]. Although the chiral magnetoelectric effect could similarly give rise to a magnetic dipole excitation, it vanishes due to the symmetry of the nanocone, as discussed below.
@@
We first consider the metasurface in free space with meta-atoms in a square periodic lattice (Fig. 1(a).) We analyze the metasurface response using full-wave simulations and extract the collective polarizabilities, which relate the induced electric and magnetic dipole moments ($\mathbf{p}$ and $\mathbf{m}$) to the incident electric and magnetic fields ($\mathbf{E}$ and $\mathbf{H}$) [4]:
$\begin{array}[]{c}\displaystyle{\bf p}=\left[\alpha_{\rm e}\overline{\overline{I}}+\alpha_{\rm ge}\overline{\overline{J}}\right]\cdot{\bf E}+\left[\alpha_{\rm te}\overline{\overline{I}}+\alpha_{\rm om}\overline{\overline{J}}\right]\cdot{\bf H},\vskip 6.0pt\\ \displaystyle{\bf m}=\left[\alpha_{\rm m}\overline{\overline{I}}+\alpha_{\rm gm}\overline{\overline{J}}\right]\cdot{\bf H}+\left[\alpha_{\rm te}\overline{\overline{I}}+\alpha_{\rm om}\overline{\overline{J}}\right]\cdot{\bf E},\end{array}$ (1)
where $\overline{\overline{I}}$ represents the two-dimensional identity matrix, and $\overline{\overline{J}}$ denotes the antisymmetric (skew-symmetric) two-dimensional matrix characterized by the nonzero components $J_{yx}=-J_{xy}=1$. We limit our analysis to the dipolar moments since they provide a dominant contribution to the total scattering of the considered nanocone (see Supplementary Section 3). The quadrupole contribution is neglected as an approximation because it is at least one order of magnitude smaller than the corresponding dipolar contributions. Thus, we model the nanocones as bianisotropic meta-atoms (with weak spatial dispersion [50]) described by the local material relations (1). Light is incident on the metasurface along the $z$-axis (normal to the surface), which is also the uniaxial direction of the metasurface. The collective polarizabilities $\alpha_{\rm e}$ and $\alpha_{\rm m}$ represent the ordinary reciprocal electric and magnetic responses, respectively. The terms $\alpha_{\rm ge}$ and $\alpha_{\rm gm}$ correspond to the nonreciprocal gyroelectric and gyromagnetic responses, while $\alpha_{\rm te}$ and $\alpha_{\rm om}$ describe the nonreciprocal Tellegen and reciprocal omega responses, respectively. All the polarizabilities are defined with respect to the base of the nanocone. It should be noted that the collective polarizabilities of the metasurface in (1) depend on the angle of incidence, even when the meta-atoms are perfectly local, i.e., characterized by angle-independent individual polarizabilities. This dependence arises from the mutual interactions among the meta-atoms within the lattice, which constitute an angle-dependent phenomenon [68]. Nevertheless, under normal incidence and for sufficiently sparse metasurfaces, the collective and individual polarizabilities of a dipolar meta-atom are qualitatively similar and differ in magnitudes by a nearly constant small factor [36, 43]. Therefore, in the following, we describe the Tellegen metasurface using collective polarizabilities, as they are significantly easier to extract experimentally and have the same order of magnitude as the individual ones.
@@
By evaluating the full scattering matrix of the metasurface through full-wave simulations, we compute the three nonreciprocal collective polarizabilities (see Supplementary Section 3), as shown in Figs. 1(b)–(d). From Fig. 1(e), it can be observed that all three polarizabilities exhibit similar magnitudes, necessitating illumination from both sides for their accurate extraction. At approximately 830 nm, the Tellegen response exhibits a Lorentzian-type resonance. Additionally, the Tellegen meta-atom demonstrates a resonant gyromagnetic response, whose presence at optical frequencies was only recently observed in metamaterials [67, 65]. The origin of this gyromagnetic effect can be explained as the combination of weak spatial dispersion and the gyroelectric effect inside the nanocone (see Supplementary Section 3).
@@
II.2 Experimental extraction of the Tellegen effect
@@
where we introduce $C_{\rm ge}$, $C_{\rm gm}$, and $C_{\rm te}$ as the weight amplitudes of the gyroelectric, gyromagnetic, and Tellegen effects in the cross-polarized reflected light. These amplitudes depend on the free-space wavelength $\lambda$, periodicity of the metasurface $a$, as well as material properties and thicknesses of the dielectric spacer and metallic layer. Importantly, these coefficients are independent of the meta-atoms and represent only the environment surrounding the metasurface. The free-space impedance $\eta_{0}$ is included in (3) to ensure that the polarizabilities remain normalized and have consistent dimensions. In our analysis, we assume that substrate-induced bianisotropy effects [32, 54, 2] can be neglected, as confirmed by our full-wave simulations (see Supplementary Section 4).
@@
To determine the Tellegen effect of the proposed optical metasurface, we use the three-metasurface method illustrated in Fig. 2. Three metasurfaces with dielectric spacer thicknesses $h_{1}=0$ nm, $h_{2}=60$ nm, and $h_{3}=120$ nm are fabricated, approximately corresponding to the maxima of the weight amplitudes within the wavelength range of interest (around 830 nm). For each metasurface $(j)$, the complex cross-polarized reflection coefficient is obtained from conventional magneto-optical Kerr measurements as $R_{\mathrm{cr}}^{(j)}=(\theta_{j}+i\epsilon_{j})R_{\mathrm{co}}^{(j)}$, where $\theta_{j}$ and $\epsilon_{j}$ are the Kerr rotation and ellipticity, respectively, and $R_{\mathrm{co}}^{(j)}$ is the co-polarized complex reflection coefficient [3]. The three measured $R_{\mathrm{cr}}^{(j)}$ values, together with the corresponding calculated weight amplitudes $C_{\mathrm{ge}}$, $C_{\mathrm{gm}}$, and $C_{\mathrm{te}}$, provide a system of linearly independent equations that can be solved for the unknown polarizabilities $\alpha_{\mathrm{ge}}$, $\alpha_{\mathrm{gm}}$, and $\alpha_{\mathrm{te}}$ (see Supplementary Section 4 for details).
II.3 Experimental realization of the optical Tellegen metasurface
Figure 3: Fabricated Tellegen metasurface and its characterization. (a)–(b) SEM images of one of the fabricated Tellegen metasurfaces with dielectric spacer thickness of $h=60$ nm (top and tilted views). The inset in (b) presents a zoomed-in image of a single meta-atom, artificially colored to visualize the cobalt and silicon parts with the averaged geometrical parameters of the nanocone used in full-wave simulations: $h_{\rm Si}=20$ nm, $h_{\rm Co}=104$ nm, $D_{\rm Si}=90$ nm, and $D_{\rm Co}=84$ nm. The averaged half-apex angle of the nanocone is $85^{\circ}$, which reflects the degree of tip rounding due to the hole-mask colloidal lithography process. (c) Schematic illustration of the polar magneto-optic Kerr effect (MOKE) experimental setup. Detailed information about the setup is available in the Methods section. (d)–(f) Experimental and simulated magneto-optical Kerr spectra for the three fabricated metasurfaces under an external magnetic field fixed at $B=1.3$ T. Here, the positive-value magnetic field denotes the direction from the base of the meta-atom to its apex. The simulated spectra correspond to periodic metasurfaces with geometric parameters averaged from the fabricated samples. (g)–(i) The gyroelectric, gyromagnetic, and Tellegen polarizabilities extracted from the data in (d)–(f).